Calculate mean from the following data using short cut method:

Marks 

Less than 10

Less than 20

Less than 30

Less than 40

Less than 50

no of students

7

18

25

35

50



1
Expert's answer
2021-06-30T13:45:59-0400

Xi is mid-point, a = 25, h = 10, di = (Xi-a)/h

Marks Fi Xi Xi - a di fi×di

< 10 7 5 -20 -2 -14

< 20 18 15 -10 -1 -18

< 30 25 25 0 0 0

< 40 35 35 10 1 35

< 50 50 45 20 2 100

Sum: 135 103

Mean = 25 + 10×103/135 = 32.63.